Sales of a software company increased linearly from $120,000 in 1996 to $180,000 in 1999
A) find an equation that expresses the sales (y) in year (x) (where (x)= 0 corresponds to 1996)
and
B) estimate the sales in 2001

Answer by lkwtcnj(8)

Since x is the year and y is sales, then the slope of the equation is the change in sales over the change in years:
m = (180,000 - 120,000) / (3 - 0) (since 1996 is 0, 1999 would be 3)
m = 60,000 / 3 = 20,000
The y-intercept is found by plugging in one of the known points:
120,000 = (0 * 20,000) + b
b = 120,000
Then the equation would be y = 20,000x + 120,000
Since 2001 would correspond to x = 5, the sales y is:
y = (20,000 * 5) + 120,000 = 220,000